The opposite angles of a parallelogram are (3x + 5)° and (61 – x)°. Find the measure of four angles.

Grade:12

3 Answers

Harshit Singh
3 years ago
Dear Student

(3x + 5)° and (61 – x)° are the opposite angles of a parallelogram.
We know that the opposite angles of a parallelogram are equal.
Therefore,
(3x + 5)° = (61 – x)°
3x + x = 61° – 5°
4x = 56°
x = 56°/4
x = 14°
⇒ 3x + 5 = 3(14) + 5 = 42 + 5 = 47
61 – x = 61 – 14 = 47
The measure of angles adjacent to the given angles = 180° – 47° = 133°
Hence, the measure of four angles of the parallelogram are 47°, 133°, 47°, and 133°.

Thanks
Mannat Arora
2 years ago
a. Find the square root of 15876 by using prime factorization method.
b. Find the square root of 121 by the method of repeated subtraction..
Mannat Arora
2 years ago
27. a. Find the square root of 15876 by using prime factorization method.
b. Find the square root of 121 by the method of repeated subtraction..